# LETS Forecast: Learning Embedology for Time Series Forecasting

### v1

@@ 2.2 Learning Dynamical Systems for Forecasting
Learning dynamical systems for time series forecasting has garnered considerable interest within the research community. Many prior works builds on Koopman’s theory (Brunton et al., 2022), which represents a nonlinear system with a linear operator in an infinite-dimensional space. Examples includes Koopman Autoencoder (Lusch et al., 2018; Takeishi et al., 2017) and K-Forecast (Lange et al., 2021). Both approximate the Koopman operator in a high-dimensional space to effectively model nonlinear dynamics. These approaches enable scalable forecasting for complex systems by simultaneously learning the measurement function and the Koopman operator.
@@ 2.3 Empirical Dynamical Modeling
EDM (Chang et al., 2017; Sugihara & May, 1990) presents an approach to model nonlinear dynamics that is different from the aforementioned works. Rooted in Takens’ theorem (Takens, 1981b), EDM relies on delay-coordinate embeddings to reconstruct the underlying attractor, thereby preserving the essential topological properties of the original dynamical system. Unlike Koopman, which linearizes nonlinear dynamics in a carefully chosen high dimensional space (approximation to an infinite-dimensional space), EDM can topologically reconstruct system dynamics using low dimensional observations, or even with a scalar observation at each time step (Takens, 1981b; Sauer et al., 1991). EDM is thus particularly attractive for real-world problems with limited observation of the system states.
@@ 2.4 Chaotic Time Series Forecasting
A related research direction focuses on forecasting chaotic time series via state space reconstruction, mirroring the underlying principles of EDM. Pioneering work by Farmer & Sidorowich (1987) introduced local approximation techniques within reconstructed state spaces using delay embeddings, facilitating short-term predictions. Subsequent studies explored the application of feedforward neural networks for learning direct mappings from reconstructed phase states to future states (Karunasinghe & Liong, 2006). Recurrent neural networks, particularly Echo State Networks (ESNs) (Jaeger & Haas, 2004), have also shown promise, with adaptations like robust ESNs (Li et al., 2012) addressing the inherent sensitivity of chaotic signals to noise and outliers. However, a significant gap remains: the development of a fully differentiable, end-to-end trainable neural network architecture that seamlessly integrates dynamical systems theory with deep learning methodologies.
@@ 3.1 Preliminaries: Takens’ Theorem and EDM
[…]
@@
[…]
@@
where the weight is given by $w_{i}=k\left(\hat{y}_{T},\hat{y}_{N_{i}}\right)$. Again, $N_{i}$ indexes the $22d+$ 2 nearest neighbors of $\hat{y}_{t}$, and $y_{N_{i}+\Delta t}$ denotes the observed data $\Delta t$ steps after $N_{i}$. We note that Equation 3 can be viewed as the Nadaraya-Watson estimator using $22d+$ 2 nearest neighbors, where the regressor connects the input of $\hat{y}_{N_{i}}$ to its output of $y_{N_{i}+\Delta t}$.
@@ 3.2 Our Approach: DeepEDM
This initial prediction allows us to concatenate the lookback window $y_{1:T}$ and the predicted window $y^{p}_{T+1:T+H}$, forming a new time series $[y_{1:T},y^{p}_{T+1:T+H}]$. DeepEDM will now operate on this extended sequence and further refine the initial prediction, bypassing EDM’s constraint on the forecasting horizon. This is particularly helpful for long-term forecasting, where $T$ might be smaller than $H$.
Time delay and encoding. DeepEDM further time-delays the extended sequence $[y_{1:T},y^{p}_{T+1:T+H}]$, and considers a learned encoder $\text{Enc}(\cdot)$ to project the time-delayed signals into a latent space. Formally, this is given by
@@ 3.3 Model Instantiation
The encoder $\text{Enc}(\cdot)$ is instantiated as a single linear operator shared across all variates. It operates on the time-delayed sequence $\hat{y}_{1:T+H}\in\mathbb{R}^{D\times\delta_{T}\times(T+H)}$, where each delay vector of dimension $\delta_{T}$ is linearly projected to a learnable latent with dimension $M\gg\delta_{T}$, resulting in the embeddings $z_{1:T+H}\in\mathbb{R}^{D\times M\times(T+H)}$ This lightweight $\text{Enc}(\cdot)$ seeks to preserve the local geometry of the time-delay embedding while enabling expressive comparisons in the latent space used for kernel regression (Eq. 6), which generates the time-delayed prediction $\bar{y}_{T+1:T+H}$.
@@ 4.3 Further Analyses
Table 3: Model design ablation. We evaluate the effects of progressively incorporating key components into our model, with metrics averaged over four prediction lengths and three random seeds (±$\pm\sigma$ shown). Each successive addition yields consistent improvements across most metrics relative to the preceding configuration.
@@
Setup and datasets. In addition to the standard temporal train-test split, we also partition the time series (variates) in ETT, Exchange, and Weather into disjoint training and testing sets, ensuring no overlap in sequence identity. For the ETT datasets, which contain 7 sequences, we train on sequences 02020–20 – 2 using only timesteps from the standard training split, and test on sequences 46464–64 – 6 using the standard test split. This $3:3$ split is necessary as several baseline models are unable to handle differing input dimensions between training and testing. Similarly, for the Exchange dataset (8 sequences), we train on the first $4$ sequences and test on the last 4. For the Weather dataset ($21$ sequences), we train on sequences 09090–90 – 9 and test on sequences 10–19.
@@ A.7 Additional Details on Synthetic Data Experiments
To capture distinct dynamical regimes of the Lorenz system, we generate two configurations: (i) Chaotic behavior: $\sigma=10.0$, $\rho=28.0$, $\beta=2.667$ and initial conditions: ($0.0,1.0,1.05$) (ii) Non-chaotic behavior: $\sigma=10.0$, $\rho=9$, $\beta=2.667$, and initial conditions: ($10.0,10.0,10.0$). For the Rössler system, we generate one chaotic configuration using $a=0.2,b=0.2,c=5.7$ and initial conditions: ($,1.,1$ 1 .).
The initial conditions were selected to ensure trajectories remain well-defined and not excessively perturbed even under the highest noise settings. Further to systematically evaluate model performance in presence of noise, we simulate noisy conditions by introducing Gaussian noise $N(0,\sigma_{noise}^{2})$, with $\sigma_{noise}\in\{0.0,0.5,1.0,1.5,2.0,2.5\}$ with higher $\sigma_{noise}$ denoting higher levels of noise. This results in a total of 18 synthetic datasets ($3$ systems $\times$ $6$ noise levels). The underlying dynamical systems and noise levels are illustrated in Figure 5.
@@
Chaotic datasets: Chaotic regimes pose significant challenges for long-term forecasting due to their inherent complexity, leading to relatively high errors. Despite these challenges, DeepEDM handles forecasting far more effectively than the baseline methods. At $\sigma=2.5$ and $H=48$, its MSE ($17.267$) is $40\%$ lower than Koopa ($28.804$), $45\%$ below iTransformer ($31.599$), and $60\%$ below Simplex ($43.548$). This advantage is also evident in the no-noise regime ($\sigma_{noise}=0$, $H=48$), where DeepEDM achieves an MSE of $10.467$—outperforming Koopa ($18.978$) by $44.85\%$, iTransformer ($18.531$) by $43.52\%$, and Simplex ($30.985$) by $66.22\%$. DeepEDM also demonstrates consistent superiority on the Rössler system, particularly in terms of MAE, further reinforcing its robustness against noise and its ability to model complex nonlinear dynamics.
@@
ECL 48 $0.1980{\scriptstyle\pm 0.0001}$ $0.2704{\scriptstyle\pm 0.0004}$ $0.1889{\scriptstyle\pm 0.0000}$ $0.2673{\scriptstyle\pm 0.0001}$ $\textbf{0.1591}{\scriptstyle\pm 0.0001}$ $\textbf{0.2463}{\scriptstyle\pm 0.0002}$ $0.1607{\scriptstyle\pm 0.0011}$ $0.2470{\scriptstyle\pm 0.0009}$
@@
ETTh1 48 $0.3412{\scriptstyle\pm 0.0012}$ $0.3630{\scriptstyle\pm 0.0015}$ $0.3369{\scriptstyle\pm 0.0006}$ $0.3654{\scriptstyle\pm 0.0003}$ $0.3280{\scriptstyle\pm 0.0006}$ $0.3599{\scriptstyle\pm 0.0002}$ $\textbf{0.3236}{\scriptstyle\pm 0.0012}$ $\textbf{0.3566}{\scriptstyle\pm 0.0004}$
@@
ETTh2 48 $0.2258{\scriptstyle\pm 0.0003}$ $0.2913{\scriptstyle\pm 0.0002}$ $\textbf{0.2237}{\scriptstyle\pm 0.0009}$ $0.2907{\scriptstyle\pm 0.0003}$ $0.2265{\scriptstyle\pm 0.0022}$ $0.2886{\scriptstyle\pm 0.0010}$ $0.2256{\scriptstyle\pm 0.0021}$ $\textbf{0.2875}{\scriptstyle\pm 0.0011}$
@@
ETTm1 48 $0.3045{\scriptstyle\pm 0.0006}$ $0.3346{\scriptstyle\pm 0.0003}$ $0.3045{\scriptstyle\pm 0.0004}$ $0.3346{\scriptstyle\pm 0.0004}$ $\textbf{0.2766}{\scriptstyle\pm 0.0009}$ $\textbf{0.3178}{\scriptstyle\pm 0.0004}$ $0.2768{\scriptstyle\pm 0.0007}$ $0.3184{\scriptstyle\pm 0.0007}$
96 $0.3043{\scriptstyle\pm 0.0015}$ $0.3377{\scriptstyle\pm 0.0011}$ $0.3035{\scriptstyle\pm 0.0010}$ $0.3367{\scriptstyle\pm 0.0004}$ $0.2924{\scriptstyle\pm 0.0017}$ $0.3305{\scriptstyle\pm 0.0010}$ $\textbf{0.2885}{\scriptstyle\pm 0.0003}$ $\textbf{0.3279}{\scriptstyle\pm 0.0006}$
@@
ETTm2 48 $0.1439{\scriptstyle\pm 0.0002}$ $0.2334{\scriptstyle\pm 0.0000}$ $0.1437{\scriptstyle\pm 0.0003}$ $0.2335{\scriptstyle\pm 0.0001}$ $0.1338{\scriptstyle\pm 0.0006}$ $0.2221{\scriptstyle\pm 0.0008}$ $\textbf{0.1330}{\scriptstyle\pm 0.0006}$ $\textbf{0.2210}{\scriptstyle\pm 0.0005}$
@@
144 $\textbf{0.1993}{\scriptstyle\pm 0.0004}$ $0.2708{\scriptstyle\pm 0.0001}$ $\textbf{0.1993}{\scriptstyle\pm 0.0009}$ $\textbf{0.2695}{\scriptstyle\pm 0.0003}$ $0.2011{\scriptstyle\pm 0.0042}$ $0.2711{\scriptstyle\pm 0.0027}$ $0.2020{\scriptstyle\pm 0.0042}$ $0.2714{\scriptstyle\pm 0.0025}$
@@
Traffic 48 $0.6978{\scriptstyle\pm 0.0005}$ $0.4131{\scriptstyle\pm 0.0004}$ $0.5621{\scriptstyle\pm 0.0012}$ $0.3636{\scriptstyle\pm 0.0003}$ $\textbf{0.4374}{\scriptstyle\pm 0.0053}$ $0.2863{\scriptstyle\pm 0.0013}$ $0.4501{\scriptstyle\pm 0.0008}$ $\textbf{0.2859}{\scriptstyle\pm 0.0011}$
@@
Exchange 48 $0.0430{\scriptstyle\pm 0.0003}$ $0.1428{\scriptstyle\pm 0.0006}$ $0.0423{\scriptstyle\pm 0.0001}$ $\textbf{0.1409}{\scriptstyle\pm 0.0000}$ $0.0427{\scriptstyle\pm 0.0003}$ $0.1422{\scriptstyle\pm 0.0007}$ $\textbf{0.0422}{\scriptstyle\pm 0.0003}$ $0.1411{\scriptstyle\pm 0.0006}$
96 $\textbf{0.0841}{\scriptstyle\pm 0.0004}$ $\textbf{0.2018}{\scriptstyle\pm 0.0006}$ $0.0853{\scriptstyle\pm 0.0018}$ $0.2029{\scriptstyle\pm 0.0021}$ $0.0853{\scriptstyle\pm 0.0014}$ $0.2037{\scriptstyle\pm 0.0010}$ $0.0870{\scriptstyle\pm 0.0024}$ $0.2044{\scriptstyle\pm 0.0023}$
@@
ILI 24 $2.2266{\scriptstyle\pm 0.0111}$ $0.9375{\scriptstyle\pm 0.0049}$ $2.2114{\scriptstyle\pm 0.0117}$ $0.8901{\scriptstyle\pm 0.0066}$ $1.8514{\scriptstyle\pm 0.1304}$ $0.8193{\scriptstyle\pm 0.0150}$ $\textbf{1.7489}{\scriptstyle\pm 0.1807}$ $\textbf{0.7896}{\scriptstyle\pm 0.0362}$
@@
Weather 48 $0.1665{\scriptstyle\pm 0.0012}$ $0.1919{\scriptstyle\pm 0.0029}$ $0.1618{\scriptstyle\pm 0.0012}$ $0.1870{\scriptstyle\pm 0.0013}$ $0.1396{\scriptstyle\pm 0.0020}$ $0.1705{\scriptstyle\pm 0.0028}$ $\textbf{0.1382}{\scriptstyle\pm 0.0010}$ $\textbf{0.1683}{\scriptstyle\pm 0.0013}$

### v2

@@ 2.2 Learning Dynamical Systems for Forecasting
Learning dynamical systems for time series forecasting has garnered considerable interest within the research community. Many prior works builds on Koopman’s theory (Mezić, 2021; Brunton et al., 2022), which represents a nonlinear system with a linear operator in an infinite-dimensional space. Examples includes Koopman Autoencoder (Lusch et al., 2018; Takeishi et al., 2017) and K-Forecast (Lange et al., 2021). Both approximate the Koopman operator in a high-dimensional space to effectively model nonlinear dynamics. These approaches enable scalable forecasting for complex systems by simultaneously learning the measurement function and the Koopman operator.
@@ 2.3 Empirical Dynamical Modeling
EDM (Chang et al., 2017; Sugihara & May, 1990) presents an alternative approach to model nonlinear dynamics. Rooted in Takens’ theorem (Takens, 1981b), it relies on delay-coordinate embeddings to reconstruct the underlying attractor, thereby preserving the essential topological properties of the original dynamical system. Unlike Koopman, which linearizes nonlinear dynamics in a carefully chosen high dimensional space (approximation to an infinite-dimensional space), EDM can topologically reconstruct system dynamics using low dimensional observations, or even with a scalar observation at each time step (Takens, 1981b; Sauer et al., 1991). EDM is thus particularly attractive for problems with limited observation of the system states.
A line of prior research (Mezić & Banaszuk, 2004; Arbabi & Mezić, 2017; Mezić, 2022) has also explored the connection between Takens’ theory and Koopman operator theory. Hankel-DMD (Arbabi & Mezić, 2017) applies dynamic mode decomposition to time-delayed time series, effectively performing Koopman spectral analysis on Takens’ embeddings. Numerical issues associated with Hankel-DMD have been discussed in (Mezić, 2022).
@@ 2.4 Chaotic Time Series Forecasting
A related research direction focuses on forecasting chaotic time series via state space reconstruction, mirroring the underlying principles of EDM. Pioneering work by (Farmer & Sidorowich, 1987) introduced local approximation techniques within reconstructed state spaces using delay embeddings for short-term predictions. Subsequent studies explored feedforward neural networks for learning direct mappings from reconstructed phase states to future states (Karunasinghe & Liong, 2006). Recurrent models, especially Echo State Networks (ESNs) (Jaeger & Haas, 2004), and variants like robust ESNs (Li et al., 2012), further improved resilience to noise and outliers.
@@ 3.1 Preliminaries: Takens’ Theorem and EDM
The (2dd+1)-D vectors $\{\displaystyle\left(\bm{h}(x),\bm{h}(\bm{\Phi}(x)),...,\bm{h}(\bm{\Phi}^{2d}(x)\right)\}$ thus preserve the topology of the states $z_{t}$. By setting $x=x_{t-2d}$, it is easy to note that these vectors are $\{[y_{t-2d},y_{t-2d+1},\dots,y_{t}]\}$, i.e., a time-delayed version of the observed time series. The theorem thus states that given a time series of 1D measurement $y_{t}$, its time delayed version $\hat{y}_{1:t},\hat{y}_{t}=\left(y_{t-2d},y_{t-2d+1},...,y_{t}\right)$ has a similar topology with the states $x_{0:t}$. Therefore, we can instead model the induced dynamics of $\hat{y}_{0:t}$ to recover properties of the underlying dynamics of $x_{1:t}$, as illustrated in Figure 1(a).
@@
Simplex projection assumes univariate measurements $y_{1:T}$ and considers its time-delayed version $\hat{y}_{1:T}$ with $\hat{y}_{t}\in\mathbb{R}^{2d+1}$, i.e., time-delayed by $2d+1$ steps. To forecast a future time step $y_{T+\Delta t}$ ($\Delta t\geq 1$), it first finds $2d+2$ nearest neighbors $\{\hat{y}_{N_{i}}\},i\in[1,...,2d+2]$ for $\hat{y}_{T}$ using a pre-specific similarity metric, i.e., a kernel function $k(\hat{y},\hat{y}^{\prime})$ [^1] The original method in (Sugihara & May, 1990) used the radial basis function kernel $k(\hat{y},\hat{y}^{\prime})=\exp(-\|\hat{y}-\hat{y}^{\prime}\|^{2}/2\sigma^{2})$.. These nearest neighbors $\{\hat{y}_{N_{i}}\}$ are assumed to define a simplex in the 2dd+1-D space of $\hat{y}$, i.e., a geometric structure that generalizes a triangle (in 2D) or a tetrahedron (in 3D) to arbitrary dimensions. $y_{T+\Delta t}$ is then predicted by a linear re-weighting on this simplex, given by
@@
where the weight is given by $w_{i}=k\left(\hat{y}_{T},\hat{y}_{N_{i}}\right)$. Again, $N_{i}$ indexes the $2d+2$ nearest neighbors of $\hat{y}_{t}$, and $y_{N_{i}+\Delta t}$ denotes the observed data $\Delta t$ steps after $N_{i}$. We note that Equation 3 can be viewed as the Nadaraya-Watson estimator using $2d+2$ nearest neighbors, where the regressor connects the input of $\hat{y}_{N_{i}}$ to its output of $y_{N_{i}+\Delta t}$.
@@ 3.2 Our Approach: DeepEDM
This initial prediction allows us to concatenate the lookback window $y_{1:T}$ and the predicted window $y^{p}_{T+1:T+H}$, forming a new time series $y_{1:T},y^{p}_{T+1:T+H}]$. DeepEDM will now operate on this extended sequence and further refine the initial prediction, bypassing EDM’s constraint on the forecasting horizon. This is particularly helpful for long-term forecasting, where $T$ might be smaller than $H$. Time delay and encoding. DeepEDM further time-delays the extended sequence $y_{1:T},y^{p}_{T+1:T+H}]$, and considers a learned encoder $\text{Enc}(\cdot)$ to project the time-delayed signals into a latent space. Formally, this is given by
@@ 3.3 Model Instantiation
The encoder $\text{Enc}(\cdot)$ is instantiated as a single linear operator shared across all variates. It operates on the time-delayed sequence $\hat{y}_{1:T+H}\in\mathbb{R}^{D\times\delta_{T}\times(T+H)}$, where each delay vector of dimension $\delta_{T}$ is linearly projected to a learnable latent with dimension $M\gg\delta_{T}$, resulting in the embeddings $z_{1:T+H}\in\mathbb{R}^{D\times M\times(T+H)}$ . This lightweight $\text{Enc}(\cdot)$ seeks to preserve the local geometry of the time-delay embedding while enabling expressive comparisons in the latent space used for kernel regression (Eq. 6), which generates the time-delayed prediction $\bar{y}_{T+1:T+H}$.
@@ 4.3 Further Analyses
Table 3: Model design ablation. We evaluate the effects of progressively incorporating key components into our model, with metrics averaged over four prediction lengths and three random seeds ($\pm\sigma$ shown). Each successive addition yields consistent improvements across most metrics relative to the preceding configuration.
@@
Setup and datasets. In addition to the standard temporal train-test split, we also partition the time series (variates) in ETT, Exchange, and Weather into disjoint training and testing sets, ensuring no overlap in sequence identity. For the ETT datasets, which contain 7 sequences, we train on sequences 020–2 using only timesteps from the standard training split, and test on sequences 464–6 using the standard test split. This $3:3$ split is necessary as several baseline models are unable to handle differing input dimensions between training and testing. Similarly, for the Exchange dataset (8 sequences), we train on the first $4$ sequences and test on the last 4. For the Weather dataset ($21$ sequences), we train on sequences 090–9 and test on sequences 10–19.
@@ A.7 Additional Details on Synthetic Data Experiments
To capture distinct dynamical regimes of the Lorenz system, we generate two configurations: (i) Chaotic behavior: $\sigma=10.0$, $\rho=28.0$, $\beta=2.667$ and initial conditions: ($0.0,1.0,1.05$) (ii) Non-chaotic behavior: $\sigma=10.0$, $\rho=9$, $\beta=2.667$, and initial conditions: ($10.0,10.0,10.0$). For the Rössler system, we generate one chaotic configuration using $a=0.2,b=0.2,c=5.7$ and initial conditions: ($1.,1.,1.$).
The initial conditions were selected to ensure trajectories remain well-defined and not excessively perturbed even under the highest noise settings. Further to systematically evaluate model performance in presence of noise, we simulate noisy conditions by introducing Gaussian noise $N(0,\sigma_{noise}^{2})$, with $\sigma_{noise}\in\{0.0,0.5,1.0,1.5,2.0,2.5\}$ with higher $\sigma_{noise}$ denoting higher levels of noise. This results in a total of 18 synthetic datasets ($3$ systems $\times$ 66 noise levels). The underlying dynamical systems and noise levels are illustrated in Figure 5.
@@
Chaotic datasets: Chaotic regimes pose significant challenges for long-term forecasting due to their inherent complexity, leading to relatively high errors. Despite these challenges, DeepEDM handles forecasting far more effectively than the baseline methods. At $\sigma=2.5$ and $H=48$, its MSE ($17.267$) is $40\%$ lower than Koopa ($28.804$), $45\%$ below iTransformer ($31.599$), and $60\%$ below Simplex ($43.548$). This advantage is also evident in the no-noise regime ($\sigma_{noise}=0$, $H=48$), where DeepEDM achieves an MSE of 10.46710.467—outperforming Koopa ($18.978$) by $44.85\%$, iTransformer ($18.531$) by $43.52\%$, and Simplex ($30.985$) by $66.22\%$. DeepEDM also demonstrates consistent superiority on the Rössler system, particularly in terms of MAE, further reinforcing its robustness against noise and its ability to model complex nonlinear dynamics.
@@
ECL
48 $0.1980{\scriptstyle\pm 0.0001}$ $0.2704{\scriptstyle\pm 0.0004}$ $0.1889{\scriptstyle\pm 0.0000}$ $0.2673{\scriptstyle\pm 0.0001}$ $\textbf{0.1591}{\scriptstyle\pm 0.0001}$ $\textbf{0.2463}{\scriptstyle\pm 0.0002}$ $0.1607{\scriptstyle\pm 0.0011}$ $0.2470{\scriptstyle\pm 0.0009}$
@@
ETTh1
48 $0.3412{\scriptstyle\pm 0.0012}$ $0.3630{\scriptstyle\pm 0.0015}$ $0.3369{\scriptstyle\pm 0.0006}$ $0.3654{\scriptstyle\pm 0.0003}$ $0.3280{\scriptstyle\pm 0.0006}$ $0.3599{\scriptstyle\pm 0.0002}$ $\textbf{0.3236}{\scriptstyle\pm 0.0012}$ $\textbf{0.3566}{\scriptstyle\pm 0.0004}$
@@
ETTh2
48 $0.2258{\scriptstyle\pm 0.0003}$ $0.2913{\scriptstyle\pm 0.0002}$ $\textbf{0.2237}{\scriptstyle\pm 0.0009}$ $0.2907{\scriptstyle\pm 0.0003}$ $0.2265{\scriptstyle\pm 0.0022}$ $0.2886{\scriptstyle\pm 0.0010}$ $0.2256{\scriptstyle\pm 0.0021}$ $\textbf{0.2875}{\scriptstyle\pm 0.0011}$
@@
ETTm1
48 $0.3045{\scriptstyle\pm 0.0006}$ $0.3346{\scriptstyle\pm 0.0003}$ $0.3045{\scriptstyle\pm 0.0004}$ $0.3346{\scriptstyle\pm 0.0004}$ $\textbf{0.2766}{\scriptstyle\pm 0.0009}$ $\textbf{0.3178}{\scriptstyle\pm 0.0004}$ $0.2768{\scriptstyle\pm 0.0007}$ $0.3184{\scriptstyle\pm 0.0007}$
96 $0.3043{\scriptstyle\pm 0.0015}$ $377{\scriptstyle\pm 0.0011} 0.303$$035{\scriptstyle\pm 0.0010}$ $0.3367{\scriptstyle\pm 0.0004}$ $0.2924{\scriptstyle\pm 0.0017}$ $0.3305{\scriptstyle\pm 0.0010}$ $\textbf{0.2885}{\scriptstyle\pm 0.0003}$ $\textbf{0.3279}{\scriptstyle\pm 0.0006}$
@@
ETTm2
48 $0.1439{\scriptstyle\pm 0.0002}$ $0.2334{\scriptstyle\pm 0.0000}$ $0.1437{\scriptstyle\pm 0.0003}$ $0.2335{\scriptstyle\pm 0.0001}$ $0.1338{\scriptstyle\pm 0.0006}$ $0.2221{\scriptstyle\pm 0.0008}$ $\textbf{0.1330}{\scriptstyle\pm 0.0006}$ $\textbf{0.2210}{\scriptstyle\pm 0.0005}$
@@
144 $\textbf{0.1993}{\scriptstyle\pm 0.0004}$ $0.2708{\scriptstyle\pm 0.0001}$ $\textbf{0.1993}{\scriptstyle\pm 0.0009}$ $\textbf{0.2695}{\scriptstyle\pm 0.0003}$ $0.2011{\scriptstyle\pm 0.0042}$ $711{\scriptstyle\pm 0.0027} 0.202$$020{\scriptstyle\pm 0.0042}$ $0.2714{\scriptstyle\pm 0.0025}$
@@
Traffic
48 $0.6978{\scriptstyle\pm 0.0005}$ $0.4131{\scriptstyle\pm 0.0004}$ $0.5621{\scriptstyle\pm 0.0012}$ $0.3636{\scriptstyle\pm 0.0003}$ $\textbf{0.4374}{\scriptstyle\pm 0.0053}$ $0.2863{\scriptstyle\pm 0.0013}$ $0.4501{\scriptstyle\pm 0.0008}$ $\textbf{0.2859}{\scriptstyle\pm 0.0011}$
@@
Exchange
[…]
96 $\textbf{0.0841}{\scriptstyle\pm 0.0004}$ $\textbf{0.2018}{\scriptstyle\pm 0.0006}$ $0.0853{\scriptstyle\pm 0.0018}$ $2029{\scriptstyle\pm 0.0021} 0.0$$0853{\scriptstyle\pm 0.0014}$ $2037{\scriptstyle\pm 0.0010} 0.0$$0870{\scriptstyle\pm 0.0024}$ $0.2044{\scriptstyle\pm 0.0023}$
@@
ILI
24 $2.2266{\scriptstyle\pm 0.0111}$ $0.9375{\scriptstyle\pm 0.0049}$ $2.2114{\scriptstyle\pm 0.0117}$ $0.8901{\scriptstyle\pm 0.0066}$ $1.8514{\scriptstyle\pm 0.1304}$ $0.8193{\scriptstyle\pm 0.0150}$ $\textbf{1.7489}{\scriptstyle\pm 0.1807}$ $\textbf{0.7896}{\scriptstyle\pm 0.0362}$
@@
Weather
48 $0.1665{\scriptstyle\pm 0.0012}$ $0.1919{\scriptstyle\pm 0.0029}$ $0.1618{\scriptstyle\pm 0.0012}$ $0.1870{\scriptstyle\pm 0.0013}$ $0.1396{\scriptstyle\pm 0.0020}$ $0.1705{\scriptstyle\pm 0.0028}$ $\textbf{0.1382}{\scriptstyle\pm 0.0010}$ $\textbf{0.1683}{\scriptstyle\pm 0.0013}$